# Shortcuts and Identifiability in Concept-based Models from a Neuro-Symbolic Lens

### v1

@@ D.3 Knowledge distillation
We consider the case where samples $(\bm{\mathrm{g}},\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}}))$ are used to distill the ground-truth knowledge. Since by Assumption 3.2 the ground-truth inference layer is deterministic, we consider the following objective for distillation. Given a subset $\mathcal{G}^{K}\subseteq\mathrm{supp}(p^{*}(\bm{\mathrm{G}}))$ the original log-likelihood objective is complemented with the following cross-entropy:
@@
i.e., the $\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})$ component of the CBM $\bm{\mathrm{\omega}}$ is increased for each $\bm{\mathrm{g}}\in\mathcal{G}^{K}$. Turned to a constraint on $\bm{\mathrm{\beta}}$ this can be written as follows:
@@ D.4 Reconstruction penalty
When focusing on reconstructing the input from the bottleneck, the probability distributions $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ may not carry enough information to completely determine the input. In fact, if $\bm{\mathrm{X}}$ depends also on stylistic variables $\bm{\mathrm{S}}\in\mathbb{R}^{q}$, it becomes impossible solely from $\bm{\mathrm{G}}$ to determine the input. This means that training a decoder only passing $\bm{\mathrm{c}}\sim p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ would not convey enough information and would reduce the benefits of the reconstruction term. Marconato et al. [2023b] have studied this setting and proposed to also include additional variables $\bm{\mathrm{Z}}\in\mathbb{R}^{q}$ in the bottleneck to overcome this problem. The encoder/decoder architecture works as follows: first the input $\bm{\mathrm{x}}$ is encoded into $p(\bm{\mathrm{C}},\bm{\mathrm{Z}}\mid\bm{\mathrm{x}})$ by the model and after sampling $(\bm{\mathrm{c}},\bm{\mathrm{z}})\sim p(\bm{\mathrm{C}},\bm{\mathrm{Z}}\mid\bm{\mathrm{x}})$ the decoder $\bm{\mathrm{d}}$ reconstruct the an input image ^$\hat{}\bm{\mathrm{x}}=\bm{\mathrm{d}}(\bm{\mathrm{c}},\bm{\mathrm{z}})$. Following, under the assumption of content-style separation, i.e., both the encoder and the decoder process independently the concept and the stylistic variables, it holds that the constraint fr maps $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ given by the reconstruction penalty result in:
@@
The proof for this can be found in [Marconato et al., 2023b, Proposition 6]. Notice that, with full support over $\mathcal{G}$, the only such maps $\bm{\mathrm{\alpha}}$ must not confuse one concept vector for another, i.e., at most there are at most $|\mathcal{G}|!$ different valid $\bm{\mathrm{\alpha}}$’s that are essentially a bijection from $\mathcal{G}$ to $\mathcal{C}$.
@@ E.1 Theoretical analysis
Deep Symbolic Learning [Daniele et al., 2023]. At inference time, this method predicts a concept vector $\bm{\mathrm{c}}\in\mathcal{C}$ from a conditional probability concept distributions, that is it implements a function $\bm{\mathrm{f}}_{DSL}:\mathbb{R}^{n}\to\mathcal{C}$ where:
@@
where $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ is a learned conditional distribution on concepts. [^5] In this case, we have that $\bm{\mathrm{f}}_{DSL}(\bm{\mathrm{X}})\neq\tilde{p}(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$. From this expression, $\bm{\mathrm{\alpha}}\in\mathcal{A}$ is defined as:
@@
Notice that, the learned $\bm{\mathrm{\beta}}:\mathcal{G}\to\Delta_{\mathcal{Y}}$ consists in a look up table from concepts vectors $\bm{\mathrm{c}}\in\mathcal{C}$, thereby the giving the conditional distribution:
@@
with $\bm{\mathrm{\omega}}$ that is defined as:
@@
This form is the same of probabilistic logic methods and we discuss the validity of Assumption 3.7 along with them.
Probabilistic Logic Models. For models like DPL [Manhaeve et al., 2018], SL [Xu et al., 2018], and SPL [Ahmed et al., 2022], $\bm{\mathrm{\omega}}$ is defined as follows:
Consider $\bm{\mathrm{c}},\bm{\mathrm{c}}^{\prime}\in\mathcal{C}$ such that:
$\bm{\mathrm{\beta}}(\bm{\mathrm{c}})\neq\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})$ (76)
Then, for any choice of $\lambda\in(0,1)$ it holds:
Notice that for all $\bm{\mathrm{y}}\in\mathcal{Y}$, it holds that by the convexity of the $\max$ operator:
$\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\big{(}\lambda\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}+(1-\lambda)\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\big{)}$ $\leq\lambda\max\big{(}\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}},\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\big{)}+(1-\lambda)\max\big{(}\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}},\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\big{)}$ (79)
where the equality holds if and only if $\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}=\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}$ and $\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}=\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}$. This proves that probabilistic logic methods satisfy Assumption 3.7.
@@
Concept Bottleneck Models. We consider CBNMs implementing a linear layer as $\bm{\mathrm{\omega}}:\Delta_{\mathcal{C}}\to\Delta_{\mathcal{Y}}$, as customary [Koh et al., 2020]. We consider the general case where the linear layer consists of a matrix $\bm{\mathrm{W}}\in\mathbb{R}^{a\times b}$ with $a:=|\mathcal{G}|$ rows and $b:=|\mathcal{Y}|$ columns. In this setting, a probability distribution $p(\bm{\mathrm{C}})\in\Delta_{\mathcal{C}}$ corresponds to the input vector that is passed to the inference layer $\bm{\mathrm{\omega}}$, implemented by the linear layer $\bm{\mathrm{W}}$ and a softmax operator. We denote with $\bm{\mathrm{w}}^{y}_{c}:=\bm{\mathrm{W}}_{c,y}$, corresponding to the $(c,y)$ component of the matrix. The index $c$ (resp. $y$) also correspond to the concept vector $\bm{\mathrm{c}}\in\mathcal{C}$ (resp. $\bm{\mathrm{y}}$), associated to it, e.g., $\bm{\mathrm{c}}=(0,1)^{\top}\in\{0,1\}^{2}$ corresponds to $c=1$. Following, we just use $\bm{\mathrm{c}}$ and $\bm{\mathrm{y}}$ to denote their corresponding index of the concepts and of the labels. Provided a concept probability vector $\bm{\mathrm{p}}\in\Delta_{\mathcal{C}}$, with components $\bm{\mathrm{p}}_{\bm{\mathrm{c}}}=p(\bm{\mathrm{C}}=\bm{\mathrm{c}})$, the conditional distribution over labels is given by the softmax operator:
$p(\bm{\mathrm{y}}\mid\bm{\mathrm{c}})$ $=\bm{\mathrm{\omega}}(\bm{\mathrm{p}})_{\bm{\mathrm{y}}}$ (81)
$=\frac{\exp\big{(}\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{W}}_{\bm{\mathrm{c}},\bm{\mathrm{y}}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big{)}}{\sum_{\bm{\mathrm{y}}^{\prime}\in\mathcal{Y}}\exp\big{(}\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{W}}_{\bm{\mathrm{c}},\bm{\mathrm{y}}^{\prime}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big{)}}$ (82)
$=\frac{\exp\big{(}\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big{)}}{\sum_{\bm{\mathrm{y}}^{\prime}\in\mathcal{Y}}\exp\big{(}\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big{)}}$ (83)
@@
In general, for any random choice of weights $\bm{\mathrm{W}}$, the CBNM may not satisfy Assumption 3.7. Notice also that, because of the form of $\bm{\mathrm{\omega}}$, the CBNM cannot express deterministic $\bm{\mathrm{\beta}}\in\mathcal{B}$. We consider the particular case associated with optimal models expressing an almost deterministic conditional distribution $\bm{\mathrm{\beta}}\in{\mathcal{B}}$, where a subset of weights $\bm{\mathrm{w}}^{\bm{\mathrm{y}}}_{\bm{\mathrm{c}}}$ is extremely high or low. Notice that this is relevant for models in the context of Assumption 3.2, where $\bm{\mathrm{\beta}}^{*}\in\mathsf{Vert}(\mathcal{B})$. We can see that in order for this to happen, we need the function $\bm{\mathrm{\beta}}$ to be peaked on concept vectors $\bm{\mathrm{c}}\in\mathcal{C}$. This can happen only when the magnitude of one $\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}$ is much higher than other $\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}}$. Hence, we formulate the following condition for CBNM:
@@
Consider a CBNM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ with weights $\bm{\mathrm{W}}=\{\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}\}$ and $M>|\mathcal{Y}-1|$. We say that the CBNM is $\log(M)$-deterministic at the vertices if for all $\bm{\mathrm{c}}\in\mathcal{C}$, there exists $\bm{\mathrm{y}}\in\mathcal{Y}$ such that:
@@
$\forall\bm{\mathrm{y}}^{\prime}\neq\bm{\mathrm{y}}\neq\bm{\mathrm{y}}^{\prime\prime},\;|\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime\prime}}|\leq\log M$ (85)
To see why this is helpful, we show that this constrained version of CBNM can flexibly approximate deterministic $\bm{\mathrm{\beta}}$’s:
@@
Consider a $\log(M)$-deterministic CBNM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$. It holds:
@@
In particular, it holds that:
$\forall\bm{\mathrm{c}}\in\mathcal{C},\;\lim_{M\to+\infty}\max_{\bm{\mathrm{y}}}\bm{\mathrm{\omega}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}=1$ (87)
@@
We now make use of the fact that the CBNM is $\log M$-deterministic and use Eq. 84 to get:
$\frac{1}{1+\sum_{\bm{\mathrm{y}}^{\prime}\neq\bm{\mathrm{y}}}\exp(-\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}+\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}})}$ $\geq\frac{1}{1+\sum_{\bm{\mathrm{y}}^{\prime}\neq\bm{\mathrm{y}}}\exp-\log M}$ (90)
@@
giving overall that:
@@
The last point of Proposition E.2 shows a viable way to get peakaed label distributions from a $\log(M)$-deterministic CBNM. This limit guaranteed that such constrained CBNMs can get closer to achieve optimal likelihood. We are now ready to prove that a $\log(M)$-deterministic CBNM respect Assumption 3.7.
@@
A CBNM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ that is $\log(M)$-deterministic (Definition E.1) satisfies Assumption 3.7, i.e., for all $\lambda\in(0,1)$ and for all $\bm{\mathrm{c}}\neq\bm{\mathrm{c}}^{\prime}$ such that $\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})_{\bm{\mathrm{y}}}\neq\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}^{\prime}\})_{\bm{\mathrm{y}}}$, it holds :
@@
Let $\bm{\mathrm{y}}_{1}:=\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\})_{\bm{\mathrm{y}}}$ and $\bm{\mathrm{y}}_{2}:=\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$ and $\lambda_{1}:=\lambda$ and $\lambda_{2}:=1-\lambda$, with $\lambda\in(0,1)$ and $\bm{\mathrm{y}}_{1}\neq\bm{\mathrm{y}}_{2}$. The proof is divided in two steps.
We first check that (1) the values taken by $\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}_{i}}$, for $i\in\{1,2\},$ are lower than those taken at the extremes. The explicit expression for $\bm{\mathrm{\omega}}(\cdot)_{\bm{\mathrm{y}}_{1}}$ is given by:

### v2

@@ D.3 Knowledge distillation
We consider the case where samples $(\bm{\mathrm{g}},\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}}))$ are used to distill the ground-truth knowledge. Since by Assumption˜3.2 the ground-truth inference layer is deterministic, we consider the following objective for distillation. Letting $\mathcal{G}^{K}\subseteq\mathrm{supp}(p^{*}(\bm{\mathrm{G}}))$ be the subset of supervised values, we augment the original log-likelihood objective with the following log-likelihood term:
@@
which penalizes the target CBM, for each $\bm{\mathrm{g}}\in\mathcal{G}^{K}$, based on the $\bm{\mathrm{\beta}}^{*}(\bm{\mathrm{g}})$ component of $\bm{\mathrm{\omega}}$, i.e., the predicted probability of the ground-truth label. This can be rewritten as a constraint on $\bm{\mathrm{\beta}}$ as follows:
@@ D.4 Reconstruction penalty
When focusing on reconstructing the input from the bottleneck, the probability distributions $p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ may not carry enough information to completely determine the input $\bm{\mathrm{X}}$. In fact, if $\bm{\mathrm{X}}$ depends also on stylistic variables $\bm{\mathrm{S}}\in\mathbb{R}^{q}$, it becomes impossible to determine the input solely from $\bm{\mathrm{G}}$. This means that training a decoder only passing $\bm{\mathrm{c}}\sim p(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ would not convey enough information and would reduce the benefits of the reconstruction term. Marconato et al. [19] have studied this setting and proposed to also include additional variables $\bm{\mathrm{Z}}\in\mathbb{R}^{q}$ in the bottleneck to overcome this problem. The encoder/decoder architecture works as follows: first the input $\bm{\mathrm{x}}$ is encoded into $p(\bm{\mathrm{C}},\bm{\mathrm{Z}}\mid\bm{\mathrm{x}})$ by the model and after sampling $(\bm{\mathrm{c}},\bm{\mathrm{z}})\sim p(\bm{\mathrm{C}},\bm{\mathrm{Z}}\mid\bm{\mathrm{x}})$ the decoder $\bm{\mathrm{d}}$ reconstruct the an input image $\hat{}\bm{\mathrm{x}}=\bm{\mathrm{d}}(\bm{\mathrm{c}},\bm{\mathrm{z}})$. Following, under the assumption of content-style separation, i.e., both the encoder and the decoder process independently the concept and the stylistic variables, it holds that the constraint fr maps $\bm{\mathrm{\alpha}}\in\mathsf{Vert}(\mathcal{A})$ given by the reconstruction penalty result in:
@@
The proof for this can be found in [19, Proposition 6]. Notice that, with full support over $\mathcal{G}$, the only such maps $\bm{\mathrm{\alpha}}$ must not confuse one concept vector for another, i.e., at most there are at most $|\mathcal{G}|!$ different valid $\bm{\mathrm{\alpha}}$’s that are essentially a bijection from $\mathcal{G}$ to $\mathcal{C}$.
@@ E.1 Theoretical analysis
Probabilistic Logic Models. For models like DeepProbLog [41], the Semantic Loss [39], Semantic-Probabilistic Layers [44], $\bm{\mathrm{\omega}}$ is defined as:
Now, take $\bm{\mathrm{c}},\bm{\mathrm{c}}^{\prime}\in\mathcal{C}$ such that: $\bm{\mathrm{\beta}}(\bm{\mathrm{c}})\neq\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})$. Then, for any $\lambda\in(0,1)$ it holds that:
Notice that for all $\bm{\mathrm{y}}\in\mathcal{Y}$, by the convexity of the $\max$ operator it holds that:
$\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\big(\lambda\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}+(1-\lambda)\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\big)\leq$ (73)
$\lambda\max\big(\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}},\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\big)+(1-\lambda)\max\big(\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}},\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}\big)$ (74)
Equality holds if and only if $\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}=\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}$ and $\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}=\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\beta}}(\bm{\mathrm{c}}^{\prime})_{\bm{\mathrm{y}}}$. This proves that probabilistic logic methods satisfy Assumption˜3.8.
Deep Symbolic Learning [50]. At inference time, DSL predicts a concept vector $\bm{\mathrm{c}}\in\mathcal{C}$ from a conditional probability concept distribution, that is, it implements a function $\bm{\mathrm{f}}_{DSL}:\mathbb{R}^{n}\to\mathcal{C}$ where:
@@
and $\tilde{p}(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$ is a learned conditional distribution on concepts. (Note that $\bm{\mathrm{f}}_{DSL}(\bm{\mathrm{X}})\neq\tilde{p}(\bm{\mathrm{C}}\mid\bm{\mathrm{X}})$.) Hence, $\bm{\mathrm{\alpha}}\in\mathcal{A}$ can be defined as:
@@
Notice that the learned $\bm{\mathrm{\beta}}:\mathcal{G}\to\Delta_{\mathcal{Y}}$ consists in a look up table from concepts vectors $\bm{\mathrm{c}}\in\mathcal{C}$, thereby giving the conditional distribution:
@@
where we used:
@@
This form matches that of probabilistic logic methods in Eq.˜70, hence Assumption˜3.8 similarly applies.
@@
Concept Bottleneck Models. We consider CBNMs implementing a linear layer as $\bm{\mathrm{\omega}}:\Delta_{\mathcal{C}}\to\Delta_{\mathcal{Y}}$, as customary [3]. The linear layer consists of a matrix $\bm{\mathrm{W}}\in\mathbb{R}^{a\times b}$ with $a:=|\mathcal{G}|$ rows and $b:=|\mathcal{Y}|$ columns. In this setting, a probability distribution $p(\bm{\mathrm{C}})\in\Delta_{\mathcal{C}}$ corresponds to the input vector passed to the inference layer $\bm{\mathrm{\omega}}$, implemented by the linear layer $\bm{\mathrm{W}}$ and a softmax operator. Notice that we can equivalently consider the tensor associated with the linear layer, $\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}$, where lower-indices $\bm{\mathrm{c}}=(c_{1},\ldots,c_{k})$ are the possible values of concepts and higher-indices $\bm{\mathrm{y}}=(y_{1},\ldots,y_{\ell})$ are the possible values of the labels. The scalar index $c$ (resp. $y$) runs over concept vectors $\bm{\mathrm{c}}\in\mathcal{C}$ (resp. $\bm{\mathrm{y}}$), e.g., for two dimensional binary concepts, $c=1$ corresponds to $\bm{\mathrm{c}}=(0,1)^{\top}\in\{0,1\}^{2}$ ($\bm{\mathrm{W}}_{1,:}=\bm{\mathrm{w}}_{01}^{\bm{\mathrm{y}}}$) and $c=2$ to $\bm{\mathrm{c}}=(1,0)^{\top}$ ($\bm{\mathrm{W}}_{2,:}=\bm{\mathrm{w}}_{10}^{\bm{\mathrm{y}}}$). In the following, we make use of the tensor notation for simplicity of exposition.
Given a concept probability vector $\bm{\mathrm{p}}\in\Delta_{\mathcal{C}}$, with components $\bm{\mathrm{p}}_{\bm{\mathrm{c}}}=p(\bm{\mathrm{C}}=\bm{\mathrm{c}})$, the conditional distribution over labels is given by the softmax operator:
$p(\bm{\mathrm{y}}\mid\bm{\mathrm{c}})=\bm{\mathrm{\omega}}(\bm{\mathrm{p}})_{\bm{\mathrm{y}}}=\frac{\exp\big(\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{W}}_{\bm{\mathrm{c}},\bm{\mathrm{y}}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big)}{\sum_{\bm{\mathrm{y}}^{\prime}\in\mathcal{Y}}\exp\big(\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{W}}_{\bm{\mathrm{c}},\bm{\mathrm{y}}^{\prime}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big)}=\frac{\exp\big(\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big)}{\sum_{\bm{\mathrm{y}}^{\prime}\in\mathcal{Y}}\exp\big(\sum_{\bm{\mathrm{c}}\in\mathcal{C}}\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}}\bm{\mathrm{p}}_{\bm{\mathrm{c}}}\big)}$ (83)
A CBNM may not satisfy Assumption˜3.8 for arbitrary choices of weights $\bm{\mathrm{W}}$. Notice also that, since $\bm{\mathrm{\omega}}$ includes a softmax operator, the $\bm{\mathrm{\beta}}\in\mathsf{Vert}(\mathcal{B})$ expressed by CBNMs cannot be deterministic . [^5] In fact, attributing near $1$ probability to a label $Y$ is only possible if the weights are extremely high in magnitude. This is a limitation for all CBNM. Specifically, in the context of Assumption˜3.2, it means that they cannot reach an optimum of maximum likelihood and therefore learn deterministic maps $\bm{\mathrm{\beta}}^{*}\in\mathsf{Vert}(\mathcal{B})$ (Assumption˜3.8).
@@
To proceed, we focus on a special case: a near-optimal CBNM that can express an “almost deterministic” conditional distribution $\bm{\mathrm{\beta}}\in{\mathcal{B}}$, where a subset of weights $\bm{\mathrm{w}}^{\bm{\mathrm{y}}}_{\bm{\mathrm{c}}}$ is very high or very low. For this to happen, we need $\bm{\mathrm{\beta}}$ to be peaked on concept vectors $\bm{\mathrm{c}}\in\mathcal{C}$. This can happen only when the magnitude of one $\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}$ is much higher than other $\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}}$. Hence, we formulate the following condition for CBNM:
@@
Consider a CBNM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ with weights $\bm{\mathrm{W}}=\{\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}\}$ and let $M>|\mathcal{Y}-1|$. We say that it is $\log(M)$-deterministic if, for all $\bm{\mathrm{c}}\in\mathcal{C}$, there exists a $\bm{\mathrm{y}}\in\mathcal{Y}$ such that:
@@
$\forall\bm{\mathrm{y}}^{\prime}\neq\bm{\mathrm{y}},\forall\bm{\mathrm{y}}^{\prime\prime}\neq\bm{\mathrm{y}},\;|\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime\prime}}|\leq\log M$ (85)
This is helpful, as any $log(M)$-deterministic CBNM can flexibly approximate deterministic $\bm{\mathrm{\beta}}$’s, as we show next:
@@
Consider a $\log(M)$-deterministic CBNM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$. We have:
@@
Also:
$\forall\bm{\mathrm{c}}\in\mathcal{C},\;\lim_{M\to+\infty}\max_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\bm{\mathrm{c}})_{\bm{\mathrm{y}}}=1$ (87)
@@
We now make use of the fact that the CBNM is $\log(M)$-deterministic and use Eq.˜84 to obtain:
$\frac{1}{1+\sum_{\bm{\mathrm{y}}^{\prime}\neq\bm{\mathrm{y}}}\exp(-(\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}}-\bm{\mathrm{w}}_{\bm{\mathrm{c}}}^{\bm{\mathrm{y}}^{\prime}}))}$ $\geq\frac{1}{1+\sum_{\bm{\mathrm{y}}^{\prime}\neq\bm{\mathrm{y}}}\exp(-\log M)}$ (90)
@@
Putting everything together yields:
@@
The last point of Proposition˜E.2 shows a viable way to get peaked label distributions from a $\log(M)$-deterministic CBNM. Specifically, the limit guarantees that these CBNMs can approach an optimal likelihood. Now, we prove that a $\log(M)$-deterministic CBNM respects Assumption˜3.8.
@@
A CBNM $(\bm{\mathrm{f}},\bm{\mathrm{\omega}})\in\mathcal{F}\times\Omega$ that is $\log(M)$-deterministic (Definition˜E.1) satisfies Assumption˜3.8, i.e., for all $\lambda\in(0,1)$ and for all $\bm{\mathrm{c}}\neq\bm{\mathrm{c}}^{\prime}$ such that $\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}\})_{\bm{\mathrm{y}}}\neq\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}^{\prime}\})_{\bm{\mathrm{y}}}$, it holds :
@@
Let $\bm{\mathrm{y}}_{1}:=\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\})_{\bm{\mathrm{y}}}$ and $\bm{\mathrm{y}}_{2}:=\operatorname*{argmax}_{\bm{\mathrm{y}}\in\mathcal{Y}}\bm{\mathrm{\omega}}(\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}}$ and $\lambda_{1}:=\lambda$ and $\lambda_{2}:=1-\lambda$, with $\lambda\in(0,1)$ and $\bm{\mathrm{y}}_{1}\neq\bm{\mathrm{y}}_{2}$. The proof consists of two steps.
First, we check that the values taken by $\bm{\mathrm{\omega}}(\lambda\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{1}\}+(1-\lambda)\mathbbm{1}\!\{\bm{\mathrm{C}}=\bm{\mathrm{c}}_{2}\})_{\bm{\mathrm{y}}_{i}}$, for $i\in\{1,2\},$ are lower than those taken at the extremes. The explicit expression for $\bm{\mathrm{\omega}}(\cdot)_{\bm{\mathrm{y}}_{1}}$ is given by: